$SPAD/src/input none.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool none.output
)set message test on
)set message auto off
)clear all
--S 1 of 3
[]
--R
--R
--R   (1)  []
--R                                                             Type: List(None)
--E 1
--S 2 of 3
[] :: List Float
--R
--R
--R   (2)  []
--R                                                            Type: List(Float)
--E 2
--S 3 of 3
[]$List(NonNegativeInteger)
--R
--R
--R   (3)  []
--R                                               Type: List(NonNegativeInteger)
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4